# DiffSampling: Enhancing Diversity and Accuracy in Neural Text Generation

### v4

@@ Abstract
Despite their growing capabilities, language models still frequently reproduce content from their training data, generate repetitive text, and favor common grammatical patterns and vocabulary. A possible cause is the decoding strategy: the most common strategies either consider only the most probable tokens, which reduces output diversity, or increase the likelihood of unlikely tokens, compromising output accuracy and correctness. In this paper, we propose DiffSampling, a new decoding method that leverages a mathematical analysis of the token probability distribution to ensure the generation of contextually appropriate text. In particular, the difference between consecutive, sorted probabilities can be used to truncate incorrect tokens. In addition, we also propose two variations of the proposed method that aim to correct the subtle inconsistencies of common sampling strategies. Experiments involving four different text-generation tasks demonstrate that our approach consistently performs at least on par with the existing methods it builds upon in terms of quality, while potentially improving output diversity.
@@ 1 Introduction
To address the subtle inconsistencies in existing truncation strategies, we introduce a family of decoding strategies called DiffSampling, based on the analysis of the probability distribution of the tokens, and in particular, on the minimum discrete derivative (i.e., the largest difference between consecutive probabilities in a sorted distribution). We propose a method for excluding incorrect tokens introduced by the smoothing distribution, along with two relaxed variants designed to promote output diversity by correcting standard methods (see Figure 1). We then provide a comprehensive evaluation of them under four different tasks, namely mathematical problem-solving tasks, extreme summarization, the divergent association task [^1] A common task in creativity research that evaluates the ability to generate semantically unrelated concepts (Olson et al., 2021)., and story generation against the most common baselines, and discuss their advantages and limitations. We show that DiffSampling consistently performs at least on par with the standard methods they aim to correct, while enhancing output diversity, especially in longer-form text generation.
@@ 2.1 Language Modeling
An autoregressive language model (LM) is a probability distribution $p_{\bm{\theta}}(\mathbf{x})$ parameterized by $\bm{\theta}$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$. The probability distribution is factorized as $p_{\bm{\theta}}(\mathbf{x})=\prod_{t=1}^{T}p_{\bm{\theta}}(x_{t}|x_{1}\ldots x_{t-1})$, and the LM is usually trained to maximize the likelihood of the true distribution $p^{*}(\mathbf{x})$ for any $\mathbf{x}$ from a reference dataset (the training set). In other words, given as input $x_{1}\ldots x_{t}$, the model learns to approximate the probability of each token from $\mathcal{V}$ being $x_{t+1}$. While this makes the model immediately capable of scoring the probability of a given text, it also allows for the generation of new sentences. Given a commonly human-written prefix (also known as a prompt) $\mathbf{x}=(x_{1}\ldots x_{P})$ of length $P$, we can decode a continuation $\mathbf{\hat{x}}=(x_{P+1}\ldots x_{T+P})$ from the LM through its factorized representation introduced before. However, we must remember that the model is trained to score and not to generate sentences. A given text might have zero probability for generation purposes (e.g., the text is syntactically incorrect), but non-zero probability for ranking purposes (Hewitt et al., 2022).
@@ 2.2 Decoding Strategies
Alternatively, nucleus (or top-$p$) sampling (Holtzman et al., 2020) reduces the token space to the smallest subset of tokens with a total probability no less than $p$. To restrict the nucleus to tokens whose information content is close to the expected one given prior context, locally typical sampling (Meister et al., 2023) focuses on the tokens with negative log-probability within a certain absolute range from the conditional entropy (and a total probability no less than $p$). Finally, Hewitt et al. (2022) assert that a language model learns a mixture of the true token distribution and a smoothing distribution to avoid infinite perplexity. For de-smoothing the distribution, they propose $\epsilon$- and $\eta$-sampling, which truncate tokens with a probability smaller than a threshold set a priori or dynamically through the entropy of the distribution, respectively. This threshold can also be set according to the magnitude of the highest probability as in min-$p$ (Minh et al., 2025), or based on the logit rather than the probability distribution (Tang et al., 2025). However, such strategies do not guarantee the exclusion of the smoothing-induced tail. Contrastive decoding (Li et al., 2023) leverages the difference in likelihood between a large language model and a smaller, less capable one to prioritize tokens with sufficiently high probability under the expert model. However, it requires access to a smaller model with an identical vocabulary, which is not always available. While conceptually aligned, our method simplifies the threshold computation and provides more intuitive guarantees on the suitability of selected tokens.
@@ 3 DiffSampling
In particular, it seems plausible that $\operatorname*{arg\,min}(\Delta p(x_{t}^{[n]}))\leq D$, i.e., it either marks the point where the true distribution ends and smoothing begins to take effect, or a point within the true distribution that separates tokens with significantly higher probabilities from the rest. Indeed, due to the inner nature of smoothing, it seems unreasonable that the maximum difference is between tokens with zero probability under the true distribution, and thus only because of the smoothing distribution.
Building on this intuition, we propose DiffSampling, a family of three decoding strategies. The first one, which we call DiffSampling-cut, leverages the aforementioned property and consists of cutting the distribution tail at the right side of the minimum discrete derivative, i.e., sampling among the tokens $x_{i},i\leq\operatorname*{arg\,min}(\Delta p(x_{t}^{[n]}))$. Due to the guarantee of selecting a correct token, which prioritizes reliability over aggressiveness, this approach can be seen as an improved greedy strategy: when the model has high confidence in a single token, it degenerates into the greedy strategy; otherwise, it preserves other appropriate tokens, increasing diversity.
@@
Overall, DiffSampling can be seen as a sampling scheme governed by two parameters, i.e., the probability-mass lower bound $p_{lb}$ and the truncated probability upper bound $p_{min}$ (where DiffSampling-cut just assumes a value of $0.0$ for the first and of $1.0$ for the second), plus the additional temperature $\tau$. The full algorithm is reported in Algorithm 1.
Algorithm 1 DiffSampling
Input: probabilities $\mathsf{probs}=[p_{t}^{[1]}\ldots p_{t}^{[N]}]$, lower bound $p_{lb}$, upper bound $p_{min}$, temperature $\tau$.
$\mathsf{sorted\_probs},\mathsf{indices}=\mathtt{sort}\!\left(\mathsf{probs}\right)$
$\mathsf{fwd\_probs}=\mathsf{sorted\_probs}[1\!:]+[0.]$
$\mathsf{delta\_probs}=\mathsf{fwd\_probs}-\mathsf{sorted\_probs}$
if $p_{min}>0.0$ then
  $lim=\operatorname*{arg\,min}\!\left(\mathsf{sorted\_probs}>p_{min}\cdot\mathsf{sorted\_probs}[0]\right)-1$
  $\mathsf{delta\_probs}[:\!lim]=0.$
else
  $\mathsf{nucleus}=\mathtt{cumsum}\!\left(\mathsf{sorted\_probs}\right)<p_{lb}$
  $\mathsf{delta\_probs}[\mathsf{nucleus}]=0.$
end if
$cut\_idx=\operatorname*{arg\,min}\!\left(\mathsf{delta\_probs}\right)$
$\mathsf{sorted\_probs}[cut\_idx\!+\!1\!:]=0.$
$\mathsf{probs}=\mathtt{sort\_by\_idx}\!\left(\mathsf{sorted\_probs},\mathsf{indices}\right)$
$\mathsf{logits}=\mathtt{log}(\mathsf{probs}/\mathtt{sum}\!\left(\mathsf{probs}\right))/\tau$
$\mathsf{probs}=\mathtt{softmax}(\mathsf{logits})$
Output: $\mathsf{probs}$
@@ 4 Illustrative Example
Table 1: Token probability comparison between top-$p$, min-$p$, and our methods, showing how they avoid treating differently tokens with very similar probabilities (reported in bold). The probabilities (in percentage) are taken from SmolLM-135M-Instruct (Ben Allal et al., 2024).
@@ 5.1 Models and Baselines
In all our experiments, we start from a state-of-the-art language model and test various decoding strategies. For the math problem-solving tasks, we use the Llama2-based MetaMath-7B-V1.0 model trained with self-supervised learning on MetaMathQA (Yu et al., 2024). For extreme text summarization and story generation, we utilize the Llama-3.2-3B model (Grattafiori et al., 2024), with both original and -Instruct versions. Finally, for the divergent association task, we consider Meta-Llama-3-8B (Grattafiori et al., 2024), using both pre-trained and DPO-tuned -Instruct versions. We study the performances of our three methods: DiffSampling-cut; DiffSampling-lb with $p_{lb}=0.9$; and DiffSampling-minp with $p_{min}=0.1$. While these values are sometimes sub-optimal (see Appendix D for a full ablation study), we chose to standardize their values to match those used for the top-$p$ and min-$p$ baselines. Indeed, we compare them with a total of 5 decoding-based baselines: greedy strategy; $\eta$-sampling (with $\eta=0.0003$); locally typical sampling (with $p=0.9$); top-$p$ sampling (with $p=0.9$); and min-$p$ sampling (with $p=0.1$). While other methods, such as selective sampling (Troshin et al., 2025), contrastive decoding (Li et al., 2023), and beam search (Roark, 2001), could also be considered, we restrict our analysis to sampling-based methods to ensure a fair comparison, selecting those with similar computational costs and operational principles to our approach.
@@ 5.2 Math Problem Solving
Experimental Setup. Solving math problems provides a useful case study for our decoding strategies, as it allows us to evaluate the correctness of solutions (as the percentage of correctly solved problems) and the diversity of procedures to arrive at the result. In particular, we consider the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) test sets; the relative prompts are reported in Appendix B. To avoid resource wasting, we focus on entries with a problem and a solution of no more than $512$ tokens.
We evaluate the quality of solutions through the ratio of correctly solved problems, i.e., with pass@1. Instead, the diversity is computed according to two methods: expectation-adjusted distinct $N$-grams (EAD) (Liu et al., 2022) and sentence embedding cosine diversity (SBERT) (Hong et al., 2024), which should evaluate syntactic and semantic diversity, respectively (Kirk et al., 2024). EAD counts the number of distinct $N$-grams tokens (averaging over $N=1\ldots 5$) and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations. The SBERT metric is $1$ minus the cosine similarity between the embeddings of the sentences. While originally based on Sentence-BERT (Reimers & Gurevych, 2019), we employ the more recent all-mpnet-base-v2 to obtain the embeddings, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens. Following Kirk et al. (2024), we compute cross-input EAD and SBERT, i.e., by considering the set of all outputs produced for a specific seed. In addition, we also compute against-greedy EAD and SBERT. Given each input, we compare the output with the greedy one by calculating the average expectation-adjusted distinct $N$-grams not present in the greedy response, and $1$ minus the cosine similarity between the two outputs, respectively. Finally, for a more fine-grained analysis, Appendix E reports a few examples of generated outputs.
@@
Experimental Results. Table 2 (left side) reports the results for the GSM8K test set. The greedy strategy achieves the highest average accuracy, closely followed by DiffSampling-cut. Among the other baselines, only locally typical sampling performs comparably, while DiffSampling-lb and DiffSampling-minp do not substantially differ from top-$p$ and min-$p$ on any metric (but different $p_{lb}$ or $p_{min}$ values can significantly improve the accuracy at a very small cost in diversity; see Appendix D).
Table 2 (right side) reports the results for the MATH test set. Here, the highest accuracy is reached by DiffSampling-cut, which also improves on the greedy strategy in terms of diversity. By contrast, our other two methods offer limited improvements over the sampling-based baselines.
@@ 5.3 Extreme Summarization
Experimental Setup. Summarizing paragraphs represents another meaningful case study since the same text can be correctly outlined in different ways. To keep the resource consumption as low as possible, we consider the eXtreme Summarization (XSum) dataset (Narayan et al., 2018), which contains pairs of documents and one-sentence summaries. In particular, we use the test partition ($11334$ entries) and exclude all entries with a tokenized document longer than $768$, obtaining $9815$ entries; then, we limit our experiment to $1000$ random samples, and we use the prompt suggested by Chhabra et al. (2024) and reported in Appendix B. Again, we aim to verify whether the summaries generated with DiffSampling are both diverse and of high quality. For diversity, we consider the per-input EAD and SBERT metrics, computed over five outputs sampled from the same prompt (Kirk et al., 2024), along with the against-greedy EAD and SBERT diversity scores introduced above. For quality assessment, we use ROUGE-$1$ (R-11) (Lin, 2004), a standard metric for summarization that evaluates the ratio of $1$-grams present in both the target and generated summaries, as well as the sentence embedding cosine similarity (SIM) between the generated and target summaries. In this way, we compute both syntactic and semantic quality metrics, as a good summary might use entirely different words while still preserving the original meaning. In addition, following Su et al. (2022), we compute the coherence (COH) between the generated output and the text to summarize through the cosine similarity between their SimCSE embeddings (Gao et al., 2021).
@@
Greedy $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.70_{\pm.00}$ $19_{\pm.00} 0.0$$_{\pm.00} 0.00±.0$$_{\pm.00} 0.0$$00_{\pm.00}$ $0.20_{\pm.00}$ $0.52_{\pm.00}$ $0.79_{\pm.01}$ $19_{\pm.00} 0.0$$_{\pm.00} 0.00±.0$$_{\pm.00} 0.0$$00_{\pm.00}$
Top-$p$ $0.24_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.01}$ $0.59_{\pm.01}$ $0.33_{\pm.00}$ $0.66_{\pm.01}$ $0.30_{\pm.01}$ $0.16_{\pm.00}$ $0.35_{\pm.01}$ $0.50_{\pm.01}$ $0.73_{\pm.01}$ $0.66_{\pm.01}$ $0.76_{\pm.01}$ $0.55_{\pm.01}$
$\eta$-sampling $0.23_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.00}$ $0.64_{\pm.01}$ $0.35_{\pm.00}$ $0.70_{\pm.01}$ $0.31_{\pm.01}$ $0.15_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.76_{\pm.01}$ $0.68_{\pm.00}$ $0.80_{\pm.01}$ $0.56_{\pm.01}$
Locally typical $0.24_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.01}$ $0.59_{\pm.01}$ $0.33_{\pm.00}$ $0.66_{\pm.01}$ $0.30_{\pm.01}$ $0.16_{\pm.00}$ $0.35_{\pm.01}$ $0.50_{\pm.01}$ $0.73_{\pm.01}$ $0.66_{\pm.01}$ $0.76_{\pm.01}$ $0.55_{\pm.01}$
Min-$p$ $0.25_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.59_{\pm.01}$ $0.30_{\pm.00}$ $0.65_{\pm.01}$ $0.28_{\pm.01}$ $0.19_{\pm.00}$ $0.45_{\pm.01}$ $0.65_{\pm.01}$ $0.63_{\pm.01}$ $0.47_{\pm.01}$ $0.61_{\pm.01}$ $0.39_{\pm.01}$
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.22_{\pm.00}$ $0.47_{\pm.01}$ $0.21_{\pm.01}$ $0.20_{\pm.00}$ $0.51_{\pm.00}$ $0.77_{\pm.01}$ $0.40_{\pm.01}$ $0.21_{\pm.00}$ $0.33_{\pm.01}$ $0.19_{\pm.01}$
DiffS.-lb $0.24_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.00}$ $0.60_{\pm.01}$ $0.33_{\pm.00}$ $0.67_{\pm.01}$ $0.30_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.49_{\pm.01}$ $0.73_{\pm.01}$ $0.67_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
DiffS.-minp $0.25_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.59_{\pm.01}$ $0.30_{\pm.00}$ $0.65_{\pm.01}$ $0.28_{\pm.01}$ $0.19_{\pm.00}$ $0.45_{\pm.01}$ $0.64_{\pm.01}$ $0.63_{\pm.01}$ $0.48_{\pm.01}$ $0.61_{\pm.01}$ $0.39_{\pm.01}$

### v5

@@ Abstract
Despite their growing capabilities, language models still frequently reproduce content from their training data, generate repetitive text, and favor common grammatical patterns and vocabulary. A possible cause is the decoding strategy: the most common strategies either consider only the most probable tokens, which reduces output diversity, or increase the likelihood of unlikely tokens, compromising output accuracy and correctness. In this paper, we propose DiffSampling, a new decoding method that leverages a mathematical analysis of the token probability distribution to ensure the generation of contextually appropriate text. In particular, the difference between consecutive, sorted probabilities can be used to truncate incorrect tokens. In addition, we also propose two variations of the proposed method that aim to correct the subtle inconsistencies of common sampling strategies. Experiments involving four different text-generation tasks demonstrate that our approach consistently performs at least on par with the existing methods it builds upon in terms of quality, despite sampling from a larger set of tokens.
@@ 1 Introduction
To address the subtle inconsistencies in existing truncation strategies, we introduce a family of decoding strategies called DiffSampling, based on the analysis of the probability distribution of the tokens, and in particular, on the minimum discrete derivative (i.e., the largest difference between consecutive probabilities in a sorted distribution). We propose a method for excluding incorrect tokens introduced by the smoothing distribution, along with two relaxed variants designed to promote output diversity by correcting standard methods (see Figure 1). We then provide a comprehensive evaluation of them under four different tasks [^1] The code and results are available at: https://github.com/giorgiofranceschelli/DiffSampling-tmlr, namely mathematical problem-solving tasks, extreme summarization, the divergent association task [^2] A common task in creativity research that evaluates the ability to generate semantically unrelated concepts (Olson et al., 2021)., and story generation against the most common baselines, and discuss their advantages and limitations. We show that DiffSampling consistently performs at least on par with the standard methods they aim to correct, while enhancing output diversity, especially in longer-form text generation.
@@ 2.1 Language Modeling
An autoregressive language model (LM) is a probability distribution $p_{\bm{\theta}}(\mathbf{x})$ parameterized by $\bm{\theta}$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$. The probability distribution is factorized as $p_{\bm{\theta}}(\mathbf{x})=\prod_{t=1}^{T}p_{\bm{\theta}}(x_{t}|x_{1}\ldots x_{t-1})$, and the LM is usually trained to maximize the likelihood of the true distribution $p_{\star}(\mathbf{x})$ for any $\mathbf{x}$ from a reference dataset (the training set). In other words, given as input $x_{1}\ldots x_{t}$, the model learns to approximate the probability of each token from $\mathcal{V}$ being $x_{t+1}$. While this makes the model immediately capable of scoring the probability of a given text, it also allows for the generation of new sentences. Given a commonly human-written prefix (also known as a prompt) $\mathbf{x}=(x_{1}\ldots x_{P})$ of length $P$, we can decode a continuation $\mathbf{\hat{x}}=(x_{P+1}\ldots x_{T+P})$ from the LM through its factorized representation introduced before. However, we must remember that the model is trained to score and not to generate sentences. A given text might have zero probability for generation purposes (e.g., the text is syntactically incorrect), but non-zero probability for ranking purposes (Hewitt et al., 2022).
@@ 2.2 Decoding Strategies
Alternatively, nucleus (or top-$p$) sampling (Holtzman et al., 2020) reduces the token space to the smallest subset of tokens with a total probability no less than $p$. To restrict the nucleus to tokens whose information content is close to the expected one given prior context, locally typical sampling (Meister et al., 2023) focuses on the tokens with negative log-probability within a certain absolute range from the conditional entropy (and a total probability no less than $p$). Finally, Hewitt et al. (2022) assert that a language model learns a mixture of the true token distribution and a smoothing distribution to avoid infinite perplexity. For de-smoothing the distribution, they propose $\epsilon$- and $\eta$-sampling, which truncate tokens with a probability smaller than a threshold set a priori or dynamically through the entropy of the distribution, respectively. This threshold can also be set according to the magnitude of the highest probability, as in min-$p$ (Minh et al., 2025), or based on the logit rather than the probability distribution (Tang et al., 2025). However, such strategies do not guarantee the exclusion of the smoothing-induced tail. Contrastive decoding (Li et al., 2023) leverages the difference in likelihood between a large language model and a smaller, less capable one to prioritize tokens with sufficiently high probability under the expert model. However, it requires access to a smaller model with an identical vocabulary, which is not always available. While conceptually aligned, our method simplifies the threshold computation and provides more intuitive guarantees on the suitability of selected tokens.
@@ 3 DiffSampling
In particular, it seems plausible that $\operatorname*{arg\,min}(\Delta p(x_{t}^{[n]}))\leq D$, i.e., it either marks the point where the true distribution ends and smoothing begins to take effect, or a point within the true distribution that separates tokens with significantly higher probabilities from the rest. Indeed, due to the inner nature of smoothing, it seems unreasonable that the maximum difference is between tokens with zero probability under the true distribution, and thus only because of the smoothing distribution (see Appendix A for a formal analysis on when $\operatorname*{arg\,min}(\Delta p(x_{t}^{[n]}))$ is provably $\leq D$).
Algorithm 1 DiffSampling
Input: probabilities $\mathsf{probs}=[p_{t}^{[1]}\ldots p_{t}^{[N]}]$, lower bound $\mathsf{p\_lb}=p_{lb}$, upper bound $\mathsf{p\_min}=p_{min}$, temperature $\mathsf{tau}=\tau$.
$\mathsf{sorted\_probs},\mathsf{indices}=\text{{sort}}\!\left(\mathsf{probs}\right)$
$\mathsf{fwd\_probs}=\mathsf{sorted\_probs}[1\!:]+[0.0]$
$\mathsf{delta\_probs}=\mathsf{fwd\_probs}-\mathsf{sorted\_probs}$
if $\mathsf{p\_min}>0.0$ then
  $\mathsf{lim}=\text{{argmin}}\!\left(\mathsf{sorted\_probs}>\mathsf{p\_min}\cdot\mathsf{sorted\_probs}[0]\right)-1$
  $\mathsf{delta\_probs}[:\!\mathsf{lim}]=0.0$
else
  $\mathsf{nucleus}=\text{{cumsum}}\!\left(\mathsf{sorted\_probs}\right)<\mathsf{p\_lb}$
  $\mathsf{delta\_probs}[\mathsf{nucleus}]=0.0$
end if
$\mathsf{cut\_idx}=\text{{argmin}}\!\left(\mathsf{delta\_probs}\right)$
$\mathsf{sorted\_probs}[\mathsf{cut\_idx}\!+\!1\!:]=0.0$
$\mathsf{probs}=\text{{sort\_by\_idx}}\!\left(\mathsf{sorted\_probs},\mathsf{indices}\right)$
$\mathsf{logits}=\text{{log}}(\mathsf{probs}/\text{{sum}}\!\left(\mathsf{probs}\right))/\mathsf{tau}$
$\mathsf{probs}=\text{{softmax}}(\mathsf{logits})$
Output: $\mathsf{probs}$.
Building on this intuition, we propose DiffSampling, a family of three decoding strategies (the full algorithm is reported in Algorithm 1). The first one, which we call DiffSampling-cut, leverages the aforementioned property and consists of cutting the distribution tail at the right side of the minimum discrete derivative, i.e., sampling among the tokens $x_{i},i\leq\operatorname*{arg\,min}(\Delta p(x_{t}^{[n]}))$. Due to the guarantee of selecting a correct token, which prioritizes reliability over aggressiveness, this approach can be seen as an improved greedy strategy: when the model has high confidence in a single token, it degenerates into the greedy strategy; otherwise, it preserves other appropriate tokens, increasing diversity. The next section provides a toy example to showcase this relation.
@@
Overall, DiffSampling can be seen as a sampling scheme governed by two parameters, i.e., the probability-mass lower bound $p_{lb}$ and the truncated probability upper bound $p_{min}$ (where DiffSampling-cut just assumes a value of $0.0$ for the first and of $1.0$ for the second), plus the additional temperature $\tau$.
@@ 4 Illustrative Example
Table 1: Token probability comparison between top-$p$, min-$p$, and our methods, showing how they avoid treating tokens with very similar probabilities differently (reported in bold). The probabilities (in percentage) are taken from SmolLM-135M-Instruct (Ben Allal et al., 2024).
@@ 5.1 Models and Baselines
In all our experiments, we start from a state-of-the-art language model and test various decoding strategies. For the math problem-solving tasks, we use the Llama2-based MetaMath-7B-V1.0 model trained with self-supervised learning on MetaMathQA (Yu et al., 2024). For extreme text summarization and story generation, we utilize the Llama-3.2-3B model (Grattafiori et al., 2024), with both original and -Instruct versions. Finally, for the divergent association task, we consider Meta-Llama-3-8B (Grattafiori et al., 2024), using both pre-trained and DPO-tuned -Instruct versions. We study the performances of our three methods: DiffSampling-cut; DiffSampling-lb with $p_{lb}=0.9$; and DiffSampling-minp with $p_{min}=0.1$. While these values are sometimes sub-optimal (see Appendix G for a full ablation study), we chose to standardize their values to match those used for the top-$p$ and min-$p$ baselines. Indeed, we compare them with a total of 5 decoding-based baselines: greedy strategy; $\eta$-sampling (with $\eta=0.0003$); locally typical sampling (with $p=0.9$); top-$p$ sampling (with $p=0.9$); and min-$p$ sampling (with $p=0.1$). While other methods, such as selective sampling (Troshin et al., 2025), contrastive decoding (Li et al., 2023), and beam search (Roark, 2001), could also be considered, we restrict our analysis to sampling-based methods to ensure a fair comparison, selecting those with similar computational costs and operational principles to our approach.
@@ 5.2 Math Problem Solving
Experimental Setup. Solving math problems provides a useful case study for our decoding strategies, as it allows us to evaluate the correctness of solutions (as the percentage of correctly solved problems) and the diversity of procedures to arrive at the result. In particular, we consider the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) test sets; the relative prompts are reported in Appendix D. To avoid resource wasting, we focus on entries with a problem and a solution of no more than $512$ tokens.
We evaluate the quality of solutions through the ratio of correctly solved problems, i.e., with pass@1. Instead, the diversity is computed according to two methods: expectation-adjusted distinct $N$-grams (EAD) (Liu et al., 2022) and sentence embedding cosine diversity (SBERT) (Hong et al., 2024), which should evaluate syntactic and semantic diversity, respectively (Kirk et al., 2024). EAD counts the number of distinct $N$-grams tokens (averaging over $N=1\ldots 5$) and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations [^3] Note that EAD is not upper-bounded. Moreover, it counts for distinct $N$-grams across all outputs, including inside the same output: the EAD of a set of equal, non-empty sentences is not 0, as each sentence will contain at least one distinct $1$-gram. In general, an EAD score cannot be considered high or low per se, but it must be compared with other EAD scores from experiments under similar conditions.. The SBERT metric is $1$ minus the cosine similarity between the embeddings of the sentences. While originally based on Sentence-BERT (Reimers and Gurevych, 2019), we employ the more recent all-mpnet-base-v2 to obtain the embeddings, as suggested by their developers [^4] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens. Following Kirk et al. (2024), we compute cross-input EAD and SBERT, i.e., by considering the set of all outputs produced for a specific seed. In addition, we also compute against-greedy EAD and SBERT. Given each input, we compare the output with the greedy one by calculating the average expectation-adjusted distinct $N$-grams not present in the greedy response, and $1$ minus the cosine similarity between the two outputs, respectively. We refer the interested reader to Appendix C for a formal definition of all used metrics. Finally, for a more fine-grained analysis, Appendix H reports a few examples of generated outputs.
@@
Experimental Results. Table 2 (left side) reports the results for the GSM8K test set. The greedy strategy achieves the highest average accuracy, closely followed by DiffSampling-cut. Among the other baselines, only locally typical sampling performs comparably, while DiffSampling-lb and DiffSampling-minp do not substantially differ from top-$p$ and min-$p$ on any metric (but different $p_{lb}$ or $p_{min}$ values can significantly improve the accuracy at a very small cost in diversity; see Appendix G).
Table 2 (right side) reports the results for the MATH test set. Here, the highest accuracy is reached by DiffSampling-cut, which also improves on the greedy strategy in terms of diversity. By contrast, our other two methods offer limited improvements over the sampling-based baselines. Notably, all methods achieve the same cross-input SBERT score, i.e., the overall diversity between all outputs is always the same across different methods, which might be due to the very similar levels of accuracy (and, therefore, due to the similar meaning of the proposed solutions).
@@ 5.3 Extreme Summarization
Experimental Setup. Summarizing paragraphs represents another meaningful case study since the same text can be correctly outlined in different ways. To keep the resource consumption as low as possible, we consider the eXtreme Summarization (XSum) dataset (Narayan et al., 2018), which contains pairs of documents and one-sentence summaries. In particular, we use the test partition ($11334$ entries) and exclude all entries with a tokenized document longer than $768$, obtaining $9815$ entries; then, we limit our experiment to $1000$ random samples, and we use the prompt suggested by Chhabra et al. (2024) and reported in Appendix D. Again, we aim to verify whether the summaries generated with DiffSampling are both diverse and of high quality. For diversity, we consider the per-input EAD and SBERT metrics, computed over five outputs sampled from the same prompt (Kirk et al., 2024), along with the against-greedy EAD and SBERT diversity scores introduced above. For quality assessment, we use ROUGE-$1$ (R-11) (Lin, 2004), a standard metric for summarization that evaluates the ratio of $1$-grams present in both the target and generated summaries, as well as the sentence embedding cosine similarity (SIM) between the generated and target summaries. In this way, we compute both syntactic and semantic quality metrics, as a good summary might use entirely different words while still preserving the original meaning. In addition, following Su et al. (2022), we compute the coherence (COH) between the generated output and the text to summarize through the cosine similarity between their SimCSE embeddings (Gao et al., 2021).
@@
Greedy $0.23_{\pm.00}$ $0.49_{\pm.01}$ $0.63_{\pm.01}$ $0.18_{\pm.00}$ - - - $0.22_{\pm.00}$ $0.51_{\pm.00}$ $0.74_{\pm.00}$ $0.19_{\pm.00}$ - - -
Top-$p$ $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.59_{\pm.01}$ $0.36_{\pm.01}$ $0.47_{\pm.01}$ $0.66_{\pm.01}$ $0.41_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.72_{\pm.01}$ $0.66_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
$\eta$-sampling $0.20_{\pm.00}$ $0.45_{\pm.01}$ $0.58_{\pm.01}$ $0.38_{\pm.01}$ $0.49_{\pm.01}$ $0.69_{\pm.01}$ $0.43_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.75_{\pm.01}$ $0.67_{\pm.00}$ $0.80_{\pm.01}$ $0.56_{\pm.01}$
Locally typical $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.59_{\pm.01}$ $0.36_{\pm.01}$ $0.47_{\pm.01}$ $0.66_{\pm.01}$ $0.41_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.72_{\pm.01}$ $0.66_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
Min-$p$ $0.22_{\pm.00}$ $0.46_{\pm.01}$ $0.61_{\pm.01}$ $0.36_{\pm.01}$ $0.43_{\pm.01}$ $0.64_{\pm.01}$ $0.38_{\pm.01}$ $0.20_{\pm.00}$ $0.44_{\pm.01}$ $0.63_{\pm.01}$ $0.65_{\pm.01}$ $0.47_{\pm.01}$ $0.62_{\pm.01}$ $0.39_{\pm.01}$
DiffS.-cut $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.45_{\pm.01}$ $0.23_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.00}$ $0.38_{\pm.01}$ $0.19_{\pm.00}$ $0.32_{\pm.01}$ $0.17_{\pm.01}$
DiffS.-lb $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.59_{\pm.01}$ $0.37_{\pm.01}$ $0.47_{\pm.01}$ $0.67_{\pm.01}$ $0.41_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.72_{\pm.01}$ $0.66_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
DiffS.-minp $0.22_{\pm.00}$ $0.46_{\pm.01}$ $0.60_{\pm.01}$ $0.35_{\pm.01}$ $0.43_{\pm.01}$ $0.64_{\pm.01}$ $0.38_{\pm.01}$ $0.20_{\pm.00}$ $0.44_{\pm.01}$ $0.62_{\pm.01}$ $0.65_{\pm.01}$ $0.47_{\pm.01}$ $0.63_{\pm.01}$ $0.39_{\pm.01}$